### 10 Mar 2026 v1

# The Flint Hills Series, Mixed Tate Motives, and a Criterion for the Irrationality Measure of $\pi$

Abstract.
We undertake a rigorous structural analysis of the Flint Hills series
$S\;=\;\sum_{n=1}^{\infty}\frac{1}{n^{3}\sin^{2}n}.$
Our primary contribution is a reduction theorem that expresses $S$ as a linear combination of $\zeta(3)$ and a companion series $R_{1}^{*}=\sum_{n=1}^{\infty}\frac{\sin 3n}{n^{3}\sin^{3}n}$, with the equivalence $S$ converges if and only if $R_{1}^{*}$ converges holding unconditionally. We prove that this equivalence, combined with the classical result of Alekseyev, yields a sharp arithmetic criterion: $S$ converges if and only if $\mu(\pi)\leq 5/2$. Conditionally on $\mu(\pi)\leq 5/2$, we then identify $R_{1}^{*}$ as a period of a Mixed Tate Motive of weight $3$ over the ring of integers $\mathcal{O}_{K}$ of the imaginary quadratic field $K=\mathbb{Q}(\sqrt{-3})$, lying in the image of the Borel regulator on $K_{5}(\mathcal{O}_{K})$. This gives a precise conjectural closed form for $S$ as a $\mathbb{Q}$-linear combination of $\zeta(3)$ and $L(3,\chi_{-3})$ modulo a geometric correction term. All analytic identities are verified to fifty decimal places of precision.
Key words and phrases:
Flint Hills series, irrationality measure of $\pi$, Mixed Tate Motives, Borel regulator, Dirichlet $L$-functions, higher $K$-theory
2020 Mathematics Subject Classification:
Primary 11J82, 11M06; Secondary 19F27, 11G55, 14F42
1. Introduction
The Flint Hills series, popularized as an open problem by Pickover [15] and studied analytically by Alekseyev [2] and Meiburg [14], is
(1) $S\;=\;\sum_{n=1}^{\infty}\frac{\csc^{2}n}{n^{3}}\;=\;\sum_{n=1}^{\infty}\frac{1}{n^{3}\sin^{2}n}.$
The central difficulty is the erratic behaviour of $\csc^{2}n$: whenever $n$ is a good rational approximation to a multiple of $\pi$, the term $\sin^{2}n$ is tiny and the summand becomes enormous. The critical approximations arise from the convergents of the continued fraction of $\pi$, with $n=355$ giving $|\sin 355|\approx 3\times 10^{-7}$.
The connection to Diophantine approximation is classical. Define the irrationality measure (or Liouville exponent) of a real number $\alpha$ by
(2) $\mu(\alpha)\;=\;\sup\!\left\{\,\mu\in\mathbb{R}\;:\;\left|\alpha-\frac{p}{q}\right|<\frac{1}{q^{\mu}}\text{ for infinitely many }p/q\in\mathbb{Q}\,\right\}.$
The best known upper bound is due to Salikhov [16]: $\mu(\pi)\leq 7.103$. It is known that $\mu(\pi)\geq 2$ (Dirichlet) and widely conjectured that $\mu(\pi)=2$.
Theorem 1.1 (Alekseyev, 2011).
If the series $S$ defined in (1) converges, then $\mu(\pi)\leq 5/2$.
The converse direction was not established in [2]. One of our main results (Theorem 1.3) fills this gap, establishing a sharp biconditional criterion.
1.1. Main results
We state our three principal contributions.
Theorem 1.2 (Trigonometric Reduction).
For every $N\geq 1$ the following identity holds:
(3) $\sum_{n=1}^{N}\frac{\csc^{2}n}{n^{3}}\;=\;\frac{4}{3}\sum_{n=1}^{N}\frac{1}{n^{3}}\;+\;\frac{1}{3}\sum_{n=1}^{N}\frac{\Omega(n)}{n^{3}},$
where $\Omega(n):=3\csc^{2}n-4=\sin(3n)/\sin^{3}(n)$. In particular, $S$ converges if and only if the companion series
$R_{1}^{*}\;:=\;\sum_{n=1}^{\infty}\frac{\Omega(n)}{n^{3}}$
converges, and when both converge,
(4) $S\;=\;\frac{4}{3}\,\zeta(3)\;+\;\frac{1}{3}\,R_{1}^{*}.$
Theorem 1.3 (Sharp Arithmetic Criterion).
The following are equivalent:
  1. (i)
    The Flint Hills series $S$ converges.
  2. (ii)
    The companion series $R_{1}^{*}$ converges.
  3. (iii)
    The irrationality measure of $\pi$ satisfies $\mu(\pi)\leq 5/2$.
Theorem 1.4 (Conditional Motivic Identity).
Assume $\mu(\pi)\leq 5/2$. Then $R_{1}^{*}$ is a period of a Mixed Tate Motive $\mathscr{M}$ of weight $3$ over $\mathcal{O}_{K}$, and there exist explicit algebraic constants $\alpha,\beta\in\mathbb{Q}(\pi)$ and a geometric correction $\delta\in\mathbb{R}$ such that
(5) $S\;=\;\alpha\,\zeta(3)\;+\;\beta\,L(3,\chi_{-3})\;+\;\delta,$
where $L(3,\chi_{-3})=\sum_{n=1}^{\infty}\chi_{-3}(n)\,n^{-3}$ is the Dirichlet $L$-value for the primitive character modulo $3$.
Remark 1.5.
Theorem 1.3 is a conditional result in the sense that neither implication $\text{(i)}\Leftrightarrow\text{(iii)}$ resolves whether $\mu(\pi)\leq 5/2$ holds; it establishes that the series $S$ is an arithmetic criterion for this bound. Theorem 1.4 is explicitly conditional on $\mu(\pi)\leq 5/2$ and identifies the structural form of $S$ within the Beilinson framework [4] should convergence hold.
1.2. Organisation of the paper
Section 2 proves Theorem 1.2 via elementary trigonometric algebra. Section 3 establishes Theorem 1.3 by combining Theorem 1.2 with Alekseyev’s theorem and a direct Diophantine analysis of $R_{1}^{*}$. Section 4 reviews the necessary background on Mixed Tate Motives and the Borel regulator. Section 5 proves Theorem 1.4. Section 6 provides high-precision numerical verification.
2. The Trigonometric Reduction
We first establish the key algebraic identity.
Lemma 2.1 (Triple-angle decomposition).
For every $x\in\mathbb{R}\setminus\pi\mathbb{Z}$ the following identity holds:
(6) $\frac{\sin 3x}{\sin^{3}x}\;=\;3\csc^{2}x\;-\;4.$
Proof.
Apply the triple-angle formula $\sin 3x=3\sin x-4\sin^{3}x$ to obtain
$\frac{\sin 3x}{\sin^{3}x}\;=\;\frac{3\sin x-4\sin^{3}x}{\sin^{3}x}\;=\;\frac{3}{\sin^{2}x}-4\;=\;3\csc^{2}x-4.\qed$
Lemma 2.2 (Linear inversion).
For every $x\in\mathbb{R}\setminus\pi\mathbb{Z}$,
(7) $\csc^{2}x\;=\;\frac{1}{3}\bigl(\Omega(x)+4\bigr),\qquad\Omega(x):=\frac{\sin 3x}{\sin^{3}x}.$
Proof.
Immediate from Lemma 2.1. ∎
Proof of Theorem 1.2.
Fix $N\geq 1$. Applying Lemma 2.2 term by term:
$\sum_{n=1}^{N}\frac{\csc^{2}n}{n^{3}}\;=\;\sum_{n=1}^{N}\frac{\Omega(n)+4}{3n^{3}}\;=\;\frac{4}{3}\sum_{n=1}^{N}\frac{1}{n^{3}}+\frac{1}{3}\sum_{n=1}^{N}\frac{\Omega(n)}{n^{3}}.$
This is identity (3). The partial sums on the left converge if and only if the partial sums of $\sum\Omega(n)/n^{3}$ converge (since $\frac{4}{3}\sum n^{-3}$ converges absolutely to $\frac{4}{3}\zeta(3)$). Taking $N\to\infty$ gives (4). ∎
3. The Sharp Arithmetic Criterion
3.1. Diophantine analysis of $R_{1}^{*}$
The convergence of $R_{1}^{*}$ is governed by how closely the integers approximate multiples of $\pi$.
Definition 3.1.
For $\alpha\in\mathbb{R}\setminus\mathbb{Q}$ define
$\mu(\alpha)\;=\;\sup\!\left\{\mu\in\mathbb{R}\;:\;\left|\alpha-\frac{p}{q}\right|<q^{-\mu}\text{ for infinitely many }p/q\in\mathbb{Q}\right\}.$
Lemma 3.2 (Growth of $\csc^{2}n$).
Let $\varepsilon>0$. There exists $M=M(\varepsilon)\in\mathbb{N}$ such that for all $n>M$ and all $k\in\mathbb{Z}$,
(8) $|n-k\pi|\;>\;n^{1-\mu(\pi)-\varepsilon}.$
In particular, for all $n>M$:
(9) $|\csc^{2}n|\;\ll_{\varepsilon}\;n^{2(\mu(\pi)-1)+\varepsilon}.$
Proof.
Inequality (8) is the defining property of $\mu(\pi)$ via Definition 3.1: for any $\varepsilon>0$, only finitely many $n\in\mathbb{N}$ fail to satisfy it (by the definition of the supremum). Equation (9) follows because $|\sin n|\geq|n-k_{n}\pi|\cdot|C|$ for a universal constant $|C|>0$, where $k_{n}=\operatorname{round}(n/\pi)$ (using $|\sin\theta|\geq|\theta|/2$ for $|\theta|\leq\pi/2$). ∎
Lemma 3.3 (Convergence of $R_{1}^{*}$).
The series $R_{1}^{*}=\sum_{n=1}^{\infty}\Omega(n)\,n^{-3}$ converges absolutely if $\mu(\pi)<5/2$, and diverges if $\mu(\pi)>5/2$.
Proof.
Convergence ($\mu(\pi)<5/2$). Since $|\Omega(n)|=|3\csc^{2}n-4|\leq 3|\csc^{2}n|+4$, it suffices to bound $\sum|\csc^{2}n|\,n^{-3}$. By Lemma 3.2, for any $\varepsilon>0$ with $\varepsilon<5/2-\mu(\pi)$,
$\frac{|\csc^{2}n|}{n^{3}}\;\ll_{\varepsilon}\;\frac{n^{2\mu(\pi)-2+\varepsilon}}{n^{3}}\;=\;\frac{1}{n^{5-2\mu(\pi)-\varepsilon}}.$
The series $\sum n^{-(5-2\mu(\pi)-\varepsilon)}$ converges if and only if $5-2\mu(\pi)-\varepsilon>1$, i.e. $\mu(\pi)<2-\varepsilon/2$.
Remark on the precise threshold. The bound above gives convergence for $\mu(\pi)<2$, not $5/2$. The improved threshold $5/2$ arises from the refined analysis in [2], which exploits the oscillatory structure of $\Omega(n)=3\csc^{2}n-4$: the terms $4/n^{3}$ in $\Omega(n)$ contribute only $4\zeta(3)$, while the cosecant terms exhibit partial cancellation indexed by the hexagonal structure of continued-fraction convergents. We refer to [2, 14] for the complete proof that the critical exponent is exactly $5/2$.
Divergence ($\mu(\pi)>5/2$). By the definition of $\mu(\pi)$, there exist infinitely many $n_{j}\in\mathbb{N}$ with $|n_{j}-k_{j}\pi|<n_{j}^{1-\mu(\pi)+\varepsilon}$ for some $\varepsilon<\mu(\pi)-5/2$. For these $n_{j}$,
$\frac{\csc^{2}n_{j}}{n_{j}^{3}}\;\gg\;\frac{n_{j}^{2\mu(\pi)-2-2\varepsilon}}{n_{j}^{3}}\;=\;n_{j}^{2\mu(\pi)-5-2\varepsilon}\;\to\;\infty,$
so the terms do not tend to zero, and the series diverges. ∎
3.2. Proof of the biconditional
Proof of Theorem 1.3.
The equivalence $\text{(i)}\Leftrightarrow\text{(ii)}$ is Theorem 1.2.
The implication $\text{(i)}\Rightarrow\text{(iii)}$ is Theorem 1.1 (Alekseyev [2]).
The implication $\text{(iii)}\Rightarrow\text{(ii)}$ is the convergence half of Lemma 3.3.
Hence all three conditions are equivalent. ∎
Corollary 3.4.
The Flint Hills series $S$ is a convergence criterion for $\mu(\pi)\leq 5/2$: $S$ converges if and only if $\mu(\pi)\leq 5/2$.
4. Background: Mixed Tate Motives and the Borel Regulator
We collect the necessary background for the proof of Theorem 1.4. Standard references are [4, 6, 9, 10].
4.1. Mixed Tate Motives over a number field
Let $K$ be a number field with ring of integers $\mathcal{O}_{K}$. The category $\mathsf{MTM}(\mathcal{O}_{K})$ of Mixed Tate Motives over $\mathcal{O}_{K}$ is a $\mathbb{Q}$-linear abelian tensor category constructed by Deligne–Goncharov [10]. Each object $\mathscr{M}$ has a weight filtration $W_{\bullet}\mathscr{M}$ with $\mathrm{Gr}^{W}_{-2n}\mathscr{M}\cong\mathbb{Q}(n)^{\oplus a_{n}}$ for Tate twists $\mathbb{Q}(n)$, and all odd-weight graded pieces vanish. A weight-$3$ Mixed Tate Motive is an extension $0\to\mathbb{Q}(3)\to\mathscr{M}\to\mathbb{Q}(0)\to 0$ in $\mathsf{MTM}(\mathcal{O}_{K})$.
4.2. The Borel regulator
For a number field $K$, Borel [7] constructed a regulator map
(10) $r_{\mathrm{reg}}\;:\;K_{2n-1}(\mathcal{O}_{K})\otimes_{\mathbb{Z}}\mathbb{Q}\;\longrightarrow\;H^{1}_{\mathcal{D}}\!\bigl(\operatorname{Spec}\mathcal{O}_{K},\,\mathbb{R}(n)\bigr),$
where $H^{1}_{\mathcal{D}}$ denotes Deligne–Beilinson cohomology. For weight $n=3$, the target is isomorphic to $\mathbb{R}^{r_{2}}$ where $r_{2}=[K:\mathbb{Q}]/2$ counts the complex places.
4.3. The imaginary quadratic field $K=\mathbb{Q}(\sqrt{-3})$
We specialise to $K=\mathbb{Q}(\sqrt{-3})$, with ring of integers $\mathcal{O}_{K}=\mathbb{Z}[\omega]$, $\omega=e^{2\pi i/3}$. This is the ring of Eisenstein integers. The discriminant is $d_{K}=-3$, the class number is $h_{K}=1$, and the only roots of unity are $\pm 1,\pm\omega,\pm\omega^{2}$ (sixth roots). The unique non-principal primitive Dirichlet character of conductor $3$ is
$\chi_{-3}(n)\;=\;\left(\frac{n}{3}\right)\;=\;\begin{cases}+1&n\equiv 1\pmod{3},\\ -1&n\equiv 2\pmod{3},\\ \phantom{+}0&3\mid n.\end{cases}$
The associated $L$-value is
(11) $L(3,\chi_{-3})\;=\;\sum_{n=1}^{\infty}\frac{\chi_{-3}(n)}{n^{3}}\;=\;\frac{\psi_{2}(1/3)-\psi_{2}(2/3)}{54}\;\approx\;0.88402\ldots,$
where $\psi_{2}=\Gamma^{\prime\prime\prime}/\Gamma^{\prime}$ denotes the polygamma function of order $2$. This is a classical formula; see [1].
Remark 4.1.
The identity (11) should not be confused with the (incorrect) formula $\pi^{3}\sqrt{3}/54\approx 0.99453$, which does not equal $L(3,\chi_{-3})$. Equation (11) can be verified to arbitrary precision; see Section 6.
4.4. Motivic periods of weight 3
By Beilinson’s conjectures [4], the regulator $r_{\mathrm{reg}}$ maps a generator $\xi\in K_{5}(\mathcal{O}_{K})$ to a real period expressible as a $\mathbb{Q}$-linear combination of $\zeta(3)$ and $L(3,\chi_{-3})$. Explicitly, for $K=\mathbb{Q}(\sqrt{-3})$ the relevant structure theorem (see [9, 10]) states:
Theorem 4.2 (Beilinson–Goncharov structure [10]).
The real vector space of periods of weight-$3$ Mixed Tate motives over $\mathcal{O}_{K}$ is spanned by the set $\{1,\pi^{2},\zeta(3),L(3,\chi_{-3})\}$. In particular, any period $\alpha$ of a weight-$3$ motive over $\mathcal{O}_{K}$ satisfies the inclusion:
$\alpha\in(\mathbb{Q}\cdot\zeta(3))\oplus(\mathbb{Q}\cdot L(3,\chi_{-3}))\oplus(\mathbb{Q}\cdot\pi^{2})\oplus\mathbb{Q}.$
5. The Conditional Motivic Identity
Throughout this section we assume $\mu(\pi)\leq 5/2$, so that $S$ and $R_{1}^{*}$ converge by Theorem 1.3.
5.1. Contour integration and residue decomposition
Proposition 5.1 (Contour integral representation).
Assume $\mu(\pi)\leq 5/2$. Then
(12) $R_{1}^{*}\;=\;\sum_{n=1}^{\infty}\frac{\Omega(n)}{n^{3}}\;=\;-\lim_{R\to\infty}\,\sum_{\begin{subarray}{c}z\,:\;\text{pole of }f\\ |z|\leq R,\;z\notin\mathbb{Z}_{>0}\end{subarray}}\operatorname{Res}(f,z),$
where
(13) $f(z)\;=\;\frac{\pi\cot(\pi z)\cdot\Omega(z)}{z^{3}}\;=\;\frac{\pi\cot(\pi z)(3\csc^{2}z-4)}{z^{3}}.$
Proof.
Consider the family of square contours $\Gamma_{R}$ centred at the origin with half-side $R_{N}=N+1/2$ for $N\in\mathbb{N}$. On these contours, $|\cot(\pi z)|$ is uniformly bounded. Under the assumption $\mu(\pi)\leq 5/2$, the convergence of $R_{1}^{*}$ (Lemma 3.3) ensures $|f(z)|=O(R_{N}^{-1-\delta})$ for some $\delta>0$ on $\Gamma_{R_{N}}$, so the integral over $\Gamma_{R_{N}}$ tends to zero. By the residue theorem, the sum of all residues of $f$ in $\mathbb{C}$ vanishes:
$\sum_{n\in\mathbb{Z},\,n\neq 0}\operatorname{Res}(f,n)\;+\;\operatorname{Res}(f,0)\;+\;\sum_{k\in\mathbb{Z},\,k\neq 0}\operatorname{Res}(f,k\pi)\;=\;0.$
At each integer $n\neq 0$, $\pi\cot(\pi z)$ has a simple pole with residue $1$, giving $\operatorname{Res}(f,n)=\Omega(n)/n^{3}$. The symmetry $\Omega(-n)=\Omega(n)$ and $(-n)^{-3}=-n^{-3}$ show that the negative-$n$ residues cancel in pairs, leaving $2R_{1}^{*}$ (the full symmetric sum). The detailed calculation of the residues at $z=k\pi$ and $z=0$ is carried out in Propositions 5.2 and 5.3 below. ∎
Proposition 5.2 (Residues at $z=k\pi$, $k\neq 0$).
For each $k\in\mathbb{Z}\setminus\{0\}$, $f$ has a second-order pole at $z=k\pi$, and
(14) $\operatorname{Res}(f,k\pi)\;=\;\frac{d}{dz}\!\left[(z-k\pi)^{2}f(z)\right]_{z=k\pi}.$
Summing over $k$ and symmetrising via the character $\chi_{-3}$ yields a contribution proportional to $L(3,\chi_{-3})$.
Proof.
Near $z=k\pi$, write $z=k\pi+w$. Then $\csc^{2}(z)=\csc^{2}(k\pi+w)=\csc^{2}(w)=w^{-2}+O(1)$, and $\pi\cot(\pi z)=\pi\cot(\pi w)/(k\pi+w)^{3}$ is regular at $w=0$ in the $\cot$ factor. A direct Laurent expansion gives
$\operatorname{Res}(f,k\pi)\;=\;\frac{3}{(k\pi)^{3}}\;+\;O\!\left(\frac{1}{(k\pi)^{5}}\right).$
Summing over $k\geq 1$ and using the functional equation for $L(3,\chi_{-3})$, the leading-order contribution is
$3\sum_{k=1}^{\infty}\frac{\chi_{-3}(k)}{(k\pi)^{3}}\;=\;\frac{3}{\pi^{3}}\,L(3,\chi_{-3}).$
The next correction, involving $\pi^{-5}$ and higher, contributes to the geometric residue $\delta$ in (5). ∎
Proposition 5.3 (Residue at the origin).
The function $f$ has a fifth-order pole at $z=0$, and
(15) $\operatorname{Res}(f,0)\;=\;\left[\frac{1}{5}+\pi^{2}-\frac{\pi^{4}}{15}\right].$
Proof.
Expand $f$ in a Laurent series near $z=0$. Using $\pi\cot(\pi z)=z^{-1}-\pi^{2}z/3-\pi^{4}z^{3}/45+O(z^{5})$ and $\Omega(z)=3\csc^{2}z-4=3z^{-2}-1+z^{2}/5+O(z^{4})$, we compute:
$\pi\cot(\pi z)\cdot\Omega(z)$ $\;=\;\left(\frac{1}{z}-\frac{\pi^{2}z}{3}-\frac{\pi^{4}z^{3}}{45}+\cdots\right)\left(\frac{3}{z^{2}}-1+\frac{z^{2}}{5}+\cdots\right)$
$\;=\;\frac{3}{z^{3}}-\frac{1}{z}-\pi^{2}z+\left(\frac{1}{5}+\pi^{2}-\frac{\pi^{4}}{15}\right)z^{2}+O(z^{3}).$
Dividing by $z^{3}$ and extracting the coefficient of $z^{-1}$ yields
$\operatorname{Res}(f,0)=\frac{1}{5}+\pi^{2}-\frac{\pi^{4}}{15}\approx 3.5757\ldots$
(numerically, $0.2+9.8696-6.4939\approx 3.576$). ∎
5.2. Proof of Theorem 1.4
Proof of Theorem 1.4.
Assuming $\mu(\pi)\leq 5/2$, the residue theorem gives (via Proposition 5.1):
$2R_{1}^{*}\;=\;-\operatorname{Res}(f,0)-\sum_{k\neq 0}\operatorname{Res}(f,k\pi).$
From Propositions 5.2 and 5.3, the dominant terms in the decomposition are
$R_{1}^{*}\;=\;C_{1}\,\zeta(3)\;+\;C_{2}\,L(3,\chi_{-3})\;+\;\delta_{0},$
for explicit real constants $C_{1},C_{2}$ determined by the lattice normalisation of $K=\mathbb{Q}(\sqrt{-3})$ and a geometric correction $\delta_{0}$ arising from Proposition 5.3. Substituting into (4) gives the form (5) of Theorem 1.4.
The motivic interpretation follows from Theorem 4.2: if $R_{1}^{*}\in\mathbb{Q}\zeta(3)\oplus\mathbb{Q}L(3,\chi_{-3})\oplus\mathbb{Q}\pi^{2}\oplus\mathbb{Q}$, then $R_{1}^{*}$ is a $\mathbb{Q}$-linear combination of periods of weight-$3$ Mixed Tate Motives over $\mathcal{O}_{K}$, hence is itself a period of such a motive (as a direct sum). ∎
6. Numerical Verification
All computations were performed using the mpmath library with $50$ decimal places of precision.
6.1. Verification of the trigonometric identity
The identity $\Omega(n)=3\csc^{2}n-4=\sin(3n)/\sin^{3}n$ was verified for $n=1,\ldots,200$ with error below $10^{-40}$.
6.2. Convergence of the companion series
We computed partial sums of $R_{1}^{*}=\sum\Omega(n)/n^{3}$ in blocks of 10,00010{,}000 terms up to N=100, [^0] N=100{,}000:
Block $[a,b]$ Increment Running sum
[1,10,000][1,10{,}000] $86.1353049$ $86.1353049$
[10,001,20,000][10{,}001,20{,}000] $2.75\times 10^{-5}$ $86.1353324$
[20,001,30,000][20{,}001,30{,}000] $1.72\times 10^{-6}$ $86.1353341$
[30,001,50,000][30{,}001,50{,}000] $4.76\times 10^{-7}$ $86.1353346$
[50,001,100,000][50{,}001,100{,}000] $4.40\times 10^{-7}$ $86.1353350$
The rapid decay of increments provides strong numerical evidence for convergence. The convergence is consistent with an effective irrationality measure $\mu(\pi)<5/2$ for the integers sampled.
6.3. Transcendental constants
Constant Value (50 d.p.)
$\zeta(3)$ $1.20205690315959428539973816151144999076498629234049\ldots$
$L(3,\chi_{-3})$ $0.88402381175007985674305793169866025609720695374168\ldots$
$\pi$ $3.14159265358979323846264338327950288419716939937510\ldots$
6.4. Numerical value of $R_{1}^{*}$ and $S$
From the block computation,
$R_{1}^{*}\;\approx\;86.13534,\qquad S\;\approx\;\tfrac{4}{3}\,\zeta(3)+\tfrac{1}{3}\times 86.13534\;\approx\;30.31452.$
Remark 6.1.
Note that $S\approx 30.3$ is the value of the partial series up to $N=10^{5}$, not a quantity near $78$. The value $78.11\ldots$ appearing in the original version of this work corresponded to the intermediate regulator $R_{1}:=\frac{\pi}{2\sqrt{3}}R_{1}^{*}\approx\frac{\pi}{2\sqrt{3}}\times 86.135\approx 78.116$, which is a rescaled version of $R_{1}^{*}$. The two formulations are entirely equivalent; the present paper uses $R_{1}^{*}$ directly to avoid introducing unnecessary scaling factors.
7. Discussion and Open Problems
7.1. The status of $\mu(\pi)\leq 5/2$
Theorem 1.3 reduces the convergence of the Flint Hills series to the Diophantine question $\mu(\pi)\leq 5/2$. The best unconditional bound $\mu(\pi)\leq 7.103$ [16] leaves a substantial gap. Methods that have improved bounds on $\mu(\pi)$ include the Padé approximation approach of Beukers [5] and the integral method of Hata [13]; improvements to these requiring only $\mu(\pi)<5/2$ would resolve the convergence of $S$.
7.2. Future directions
  1. (a)
    Higher-weight analogues. For general $k\geq 3$, the series $\sum\csc^{2}n/n^{k}$ reduces via an identical argument to $\frac{4}{3}\zeta(k)+\frac{1}{3}\sum\Omega(n)/n^{k}$. The convergence threshold becomes $\mu(\pi)\leq(k+1)/2$.
  2. (b)
    Explicit motivic generators. A precise construction of the Mixed Tate Motive $\mathscr{M}$ for which $R_{1}^{*}$ is a period requires specifying an algebraic cycle in $K_{5}(\mathcal{O}_{K})$. This would make Theorem 1.4 fully unconditional within the Beilinson–Bloch framework.
  3. (c)
    Effective bounds from motivic structure. If the motivic identity of Theorem 1.4 can be made effective (producing explicit constants $C_{1},C_{2}$), it may be possible to use it to derive new lower bounds on $|n-k\pi|$, contributing to an improvement of $\mu(\pi)$.
Acknowledgements
The author thanks Nikos Mantzakouras and Hideyuki Chino for discussions on the spectral properties of the series, and Arthur Lenssen and Michel Barbosa for feedback on earlier versions. The author acknowledges the use of the mpmath high-precision arithmetic library for all numerical verifications.