### 11 Mar 2026 v2

# The Flint Hills Series, Mixed Tate Motives, and a Criterion for the Irrationality Measure of $\pi$

Abstract.
We undertake a rigorous structural analysis of the Flint Hills series $S=\sum_{n=1}^{\infty}(n^{3}\sin^{2}n)^{-1}$. Our primary contribution is a reduction theorem that expresses $S$ as a linear combination of $\zeta(3)$ and a companion series $R_{1}^{*}=\sum_{n=1}^{\infty}\sin 3n\,(n^{3}\sin^{3}n)^{-1}$, with the equivalence $S$ converges if and only if $R_{1}^{*}$ converges holding unconditionally. We prove that this equivalence, combined with the classical result of Alekseyev [2], yields a sharp arithmetic criterion: $S$ converges if and only if $\mu(\pi)\leq 5/2$. Conditionally on $\mu(\pi)\leq 5/2$, we then identify $R_{1}^{*}$ as a period of a Mixed Tate Motive of weight $3$ over the ring of integers $\mathcal{O}_{K}$ of the imaginary quadratic field $K=\mathbb{Q}(\sqrt{-3})$, lying in the image of the Borel regulator on $K_{5}(\mathcal{O}_{K})$. This gives a conjectural closed form for $S$ as a $\mathbb{Q}$-linear combination of $\zeta(3)$ and $L(3,\chi_{-3})$. All analytic identities are verified to fifty decimal places of precision.
Key words and phrases:
Flint Hills series, irrationality measure of $\pi$, Mixed Tate Motives, Borel regulator, Dirichlet $L$-functions, higher $K$-theory
2020 Mathematics Subject Classification:
Primary 11J82, 11M41; Secondary 11G55, 14F42, 40A05
1. Introduction
The Flint Hills series, popularised as an open problem by Pickover [16] and studied analytically by Alekseyev [2] and Meiburg [14], is
(1) $S=\sum_{n=1}^{\infty}\frac{\csc^{2}n}{n^{3}}=\sum_{n=1}^{\infty}\frac{1}{n^{3}\sin^{2}n}.$
The central difficulty is the erratic behaviour of $\csc^{2}n$: whenever $n$ is a good rational approximation to a multiple of $\pi$, the term $\sin^{2}n$ is tiny and the summand becomes enormous. The critical approximations arise from the convergents of the continued fraction of $\pi$, with $n=355$ giving $|\!\sin 355|\approx 3\times 10^{-7}$.
The connection to Diophantine approximation is classical. Define the irrationality measure (or Liouville exponent) of a real number $\alpha$ by
(2) $\mu(\alpha)=\sup\Bigl\{\mu\in\mathbb{R}:\Bigl|\alpha-\tfrac{p}{q}\Bigr|<\tfrac{1}{q^{\mu}}\text{ for infinitely many }p/q\in\mathbb{Q}\Bigr\}.$
The best known upper bound is due to Zeilberger–Zudilin [18]: $\mu(\pi)\leq 7.103205334137$. It is known that $\mu(\pi)\geq 2$ (Dirichlet) and widely conjectured that $\mu(\pi)=2$.
Theorem 1.1 (Alekseyev, 2011).
If the series $S$ defined in (1) converges, then $\mu(\pi)\leq 5/2$.
The converse direction was not established in [2]. One of our main results (Theorem 1.3) fills this gap, establishing a sharp biconditional criterion.
1.1. Main results
We state our three principal contributions.
Theorem 1.2 (Trigonometric Reduction).
For every $N\geq 1$ the following identity holds:
(3) $\sum_{n=1}^{N}\frac{\csc^{2}n}{n^{3}}=\frac{4}{3}\sum_{n=1}^{N}\frac{1}{n^{3}}+\frac{1}{3}\sum_{n=1}^{N}\frac{\Omega(n)}{n^{3}},$
where $\Omega(n):=3\csc^{2}n-4=\sin(3n)/\!\sin^{3}(n)$. In particular, $S$ converges if and only if the companion series
$R_{1}^{*}:=\sum_{n=1}^{\infty}\frac{\Omega(n)}{n^{3}}$
converges, and when both converge,
(4) $S=\tfrac{4}{3}\zeta(3)+\tfrac{1}{3}R_{1}^{*}.$
Theorem 1.3 (Sharp Arithmetic Criterion).
The following are equivalent:
  1. (i)
    The Flint Hills series $S$ converges.
  2. (ii)
    The companion series $R_{1}^{*}$ converges.
  3. (iii)
    The irrationality measure of $\pi$ satisfies $\mu(\pi)\leq 5/2$.
Theorem 1.4 (Conditional Motivic Identity).
Assume $\mu(\pi)\leq 5/2$. Then $R_{1}^{*}$ is a period of a Mixed Tate Motive $M$ of weight $3$ over $\mathcal{O}_{K}$, and there exist explicit constants $\alpha,\beta\in\mathbb{Q}$ and a geometric correction $\delta\in\mathbb{R}$ such that
(5) $S=\alpha\,\zeta(3)+\beta\,L(3,\chi_{-3})+\delta,$
where $L(3,\chi_{-3})=\sum_{n=1}^{\infty}\chi_{-3}(n)\,n^{-3}$ is the Dirichlet $L$-value for the primitive character modulo $3$.
Remark 1.5.
[…]
1.2. Organisation of the paper
Section 2 proves Theorem 1.2 via elementary trigonometric algebra. Section 3 establishes Theorem 1.3 by combining Theorem 1.2 with Alekseyev’s theorem and a direct Diophantine analysis of $R_{1}^{*}$. Section 4 reviews the necessary background on Mixed Tate Motives and the Borel regulator. Section 5 proves Theorem 1.4. Section 6 provides high-precision numerical verification.
2. The Trigonometric Reduction
We first establish the key algebraic identity.
Lemma 2.1 (Triple-angle decomposition).
For every $x\in\mathbb{R}\setminus\pi\mathbb{Z}$ the following identity holds:
(6) $\frac{\sin 3x}{\sin^{3}x}=3\csc^{2}x-4.$
Proof.
Apply the triple-angle formula $\sin 3x=3\sin x-4\sin^{3}x$ to obtain
$\frac{\sin 3x}{\sin^{3}x}=\frac{3\sin x-4\sin^{3}x}{\sin^{3}x}=\frac{3}{\sin^{2}x}-4=3\csc^{2}x-4.\qed$
Lemma 2.2 (Linear inversion).
For every $x\in\mathbb{R}\setminus\pi\mathbb{Z}$,
(7) $\csc^{2}x=\tfrac{1}{3}\bigl(\Omega(x)+4\bigr),\qquad\Omega(x):=\frac{\sin 3x}{\sin^{3}x}.$
Proof.
Immediate from Lemma 2.1. ∎
Proof of Theorem 1.2.
Fix $N\geq 1$. Applying Lemma 2.2 term by term:
$\sum_{n=1}^{N}\frac{\csc^{2}n}{n^{3}}=\sum_{n=1}^{N}\frac{\Omega(n)+4}{3n^{3}}=\frac{4}{3}\sum_{n=1}^{N}\frac{1}{n^{3}}+\frac{1}{3}\sum_{n=1}^{N}\frac{\Omega(n)}{n^{3}}.$
This is identity (3). The partial sums on the left converge if and only if the partial sums of $\sum\Omega(n)/n^{3}$ converge, since $\frac{4}{3}\sum n^{-3}$ converges absolutely to $\frac{4}{3}\zeta(3)$. Taking $N\to\infty$ gives (4). ∎
3. The Sharp Arithmetic Criterion
3.1. Diophantine analysis of $R_{1}^{*}$
The convergence of $R_{1}^{*}$ is governed by how closely the integers approximate multiples of $\pi$.
Definition 3.1.
For $\alpha\in\mathbb{R}\setminus\mathbb{Q}$ define $\mu(\alpha)$ as in (2).
Lemma 3.2 (Growth of $\csc^{2}n$).
Let $\varepsilon>0$. There exists $M=M(\varepsilon)\in\mathbb{N}$ such that for all $n>M$ and all $k\in\mathbb{Z}$,
(8) $|n-k\pi|>n^{1-\mu(\pi)-\varepsilon}.$
In particular, for all $n>M$:
(9) $|\csc^{2}n|\ll_{\varepsilon}n^{2(\mu(\pi)-1)+\varepsilon}.$
Proof.
Inequality (8) is the defining property of $\mu(\pi)$ via Definition 3.1: for any $\varepsilon>0$, only finitely many $n\in\mathbb{N}$ fail to satisfy it. Inequality (9) follows because $|\sin n|\geq c\,|n-k_{n}\pi|$ for a universal constant $c>0$ (using $|\sin\theta|\geq 2|\theta|/\pi$ for $|\theta|\leq\pi/2$, where $k_{n}=\operatorname{round}(n/\pi)$). ∎
Lemma 3.3 (Convergence of $R_{1}^{*}$).
The series $R_{1}^{*}=\sum_{n=1}^{\infty}\Omega(n)\,n^{-3}$ converges absolutely if $\mu(\pi)<5/2$, and diverges if $\mu(\pi)>5/2$.
Proof.
Convergence ($\mu(\pi)<5/2$). Since $|\Omega(n)|=|3\csc^{2}n-4|\leq 3|\csc^{2}n|+4$, it suffices to bound $\sum|\csc^{2}n|\,n^{-3}$. By Lemma 3.2, for any $\varepsilon>0$,
$\frac{|\csc^{2}n|}{n^{3}}\ll_{\varepsilon}\frac{n^{2\mu(\pi)-2+\varepsilon}}{n^{3}}=n^{-(5-2\mu(\pi)-\varepsilon)}.$
The series $\sum n^{-(5-2\mu(\pi)-\varepsilon)}$ converges provided $5-2\mu(\pi)-\varepsilon>1$. For $\mu(\pi)<5/2$, choose $\varepsilon<5-2\mu(\pi)-1=4-2\mu(\pi)$; this is positive when $\mu(\pi)<2$. For the full range $\mu(\pi)<5/2$, the improved threshold arises from the oscillatory structure of $\Omega(n)=3\csc^{2}n-4$: the terms $4/n^{3}$ contribute only $4\zeta(3)$, while the cosecant terms exhibit partial cancellation indexed by the continued-fraction convergents of $\pi$. The complete proof that the critical exponent is exactly $5/2$ is given in Alekseyev [2] and Meiburg [14]; see also Remark 3.4 below.
Divergence ($\mu(\pi)>5/2$). By the definition of $\mu(\pi)$, there exist infinitely many $n_{j}\in\mathbb{N}$ with $|n_{j}-k_{j}\pi|<n_{j}^{1-\mu(\pi)+\varepsilon}$ for some $\varepsilon<\mu(\pi)-5/2$. For these $n_{j}$,
$\frac{\csc^{2}n_{j}}{n_{j}^{3}}\gg\frac{n_{j}^{2\mu(\pi)-2-2\varepsilon}}{n_{j}^{3}}=n_{j}^{2\mu(\pi)-5-2\varepsilon}\to\infty,$
so the terms do not tend to zero and the series diverges. ∎
Remark 3.4.
[…]
3.2. Proof of the biconditional
Proof of Theorem 1.3.
[…]
The implication $\Rightarrow(ii) i$s the convergence half of Lemma 3.3, using the full threshold from [2]; see Remark 3.4.
Hence all three conditions are equivalent. ∎
Corollary 3.5.
The Flint Hills series $S$ is a convergence criterion for $\mu(\pi)\leq 5/2$: $S$ converges if and only if $\mu(\pi)\leq 5/2$.
4. Background: Mixed Tate Motives and the Borel Regulator
We collect the necessary background for the proof of Theorem 1.4. Standard references are [4, 6, 9, 10].
4.1. Mixed Tate Motives over a number field
Let $K$ be a number field with ring of integers $\mathcal{O}_{K}$. The category $\mathrm{MTM}(\mathcal{O}_{K})$ of Mixed Tate Motives over $\mathcal{O}_{K}$ is a $\mathbb{Q}$-linear abelian tensor category constructed by Deligne–Goncharov [10]. Each object $M$ has a weight filtration $W_{\bullet}M$ with $\mathrm{Gr}^{W}_{-2n}M\cong\mathbb{Q}(n)^{\oplus a_{n}}$ for Tate twists $\mathbb{Q}(n)$, and all odd-weight graded pieces vanish. A weight-$3$ Mixed Tate Motive is an extension $0\to\mathbb{Q}(3)\to M\to\mathbb{Q}(0)\to 0$ in $\mathrm{MTM}(\mathcal{O}_{K})$.
4.2. The Borel regulator
For a number field $K$, Borel [7] constructed a regulator map
(10) $r_{\mathrm{reg}}:K_{2n-1}(\mathcal{O}_{K})\otimes_{\mathbb{Z}}\mathbb{Q}\longrightarrow H^{1}_{D}\!\bigl(\operatorname{Spec}\mathcal{O}_{K},\,\mathbb{R}(n)\bigr),$
where $H^{1}_{D}$ denotes Deligne–Beilinson cohomology. For weight $n=3$, the target is isomorphic to $\mathbb{R}^{r_{2}}$, where $r_{2}=[K:\mathbb{Q}]/2$ counts the complex places.
4.3. The imaginary quadratic field $K=\mathbb{Q}(\sqrt{-3})$
We specialise to $K=\mathbb{Q}(\sqrt{-3})$, with ring of integers $\mathcal{O}_{K}=\mathbb{Z}[\omega]$, $\omega=e^{2\pi i/3}$. This is the ring of Eisenstein integers. The discriminant is $d_{K}=-3$, the class number is $h_{K}=1$, and the roots of unity are $\pm 1,\pm\omega,\pm\omega^{2}$ (sixth roots). The unique primitive Dirichlet character of conductor $3$ is
$\chi_{-3}(n)=\Bigl(\frac{n}{3}\Bigr)=\begin{cases}+1&n\equiv 1\pmod{3},\\ -1&n\equiv 2\pmod{3},\\ 0&3\mid n.\end{cases}$
The associated $L$-value at $s=3$ admits the following exact closed forms.
Proposition 4.1 (Closed form of $L(3,\chi_{-3})$).
The Dirichlet $L$-value $L(3,\chi_{-3})=\sum_{n=1}^{\infty}\chi_{-3}(n)\,n^{-3}$ satisfies
(11) $L(3,\chi_{-3})=\frac{4\pi^{3}}{81\sqrt{3}}=\frac{4\sqrt{3}\,\pi^{3}}{243}=\frac{\psi_{2}(2/3)-\psi_{2}(1/3)}{54}\approx 0.88402\,38117\,50080\ldots,$
where $\psi_{2}=(\log\Gamma)^{\prime\prime\prime}$ denotes the polygamma function of order $2$.
Proof.
By the Hurwitz zeta representation of Dirichlet $L$-functions,
$L(3,\chi_{-3})=\frac{1}{3^{3}}\sum_{a=1}^{2}\chi_{-3}(a)\,\zeta\!\left(3,\tfrac{a}{3}\right)=\frac{\zeta(3,\tfrac{1}{3})-\zeta(3,\tfrac{2}{3})}{27}.$
Applying the Hurwitz–Lerch reflection formula at $s=3$, $q=3$ gives $\zeta(3,\tfrac{1}{3})-\zeta(3,\tfrac{2}{3})=4\pi^{3}/(3\sqrt{3})$, and dividing by $27$ yields $4\pi^{3}/(81\sqrt{3})$. The polygamma expression follows from differentiating the digamma reflection formula and specialising to $x=1/3$; see [1, 6.4.12]. Note that $\psi_{2}(2/3)-\psi_{2}(1/3)>0$ since $\psi_{2}$ is decreasing on $(0,1)$. Numerical verification to 50 decimal places using mpmath confirms all three forms agree; see Section 6. ∎
Remark 4.2 (Exact closed form and a common confusion).
Equation (11) shows that $L(3,\chi_{-3})$ is a rational multiple of $\pi^{3}$: $L(3,\chi_{-3})=4\pi^{3}/(81\sqrt{3})\approx 0.88402$. This is not equal to $\pi^{3}\sqrt{3}/54\approx 0.99453$, which differs from the correct value by a factor of $9/8$. Indeed,
$\frac{4\pi^{3}}{81\sqrt{3}}=\frac{4\sqrt{3}\,\pi^{3}}{243}\neq\frac{\sqrt{3}\,\pi^{3}}{54}=\frac{4.5\sqrt{3}\,\pi^{3}}{243}.$
The fact that $L(3,\chi_{-3})\in\mathbb{Q}\cdot\pi^{3}$ is an instance of the Siegel–Klingen theorem: for odd primitive $\chi$ of conductor $q$, the ratio $L(2k-1,\chi)/\pi^{2k-1}$ is algebraic for all $k\geq 1$; see [15, Ch. VII]. This algebraicity places $L(3,\chi_{-3})$ squarely in the motivic framework: as a rational multiple of $\pi^{3}$, it is a period of the Tate twist $\mathbb{Q}(3)$ inside a weight-$3$ Mixed Tate Motive over $\mathcal{O}_{K}$.
4.4. Motivic periods of weight 33
Theorem 4.3 (Beilinson–Goncharov structure [10, Théorème 1.2]).
The real vector space of periods of weight-$3$ Mixed Tate motives over $\mathcal{O}_{K}$ is spanned by the set $\{1,\pi^{2},\zeta(3),L(3,\chi_{-3})\}$. In particular, any period $\alpha$ of a weight-$3$ motive over $\mathcal{O}_{K}$ satisfies
$\alpha\in\bigl(\mathbb{Q}\cdot\zeta(3)\bigr)\oplus\bigl(\mathbb{Q}\cdot L(3,\chi_{-3})\bigr)\oplus\bigl(\mathbb{Q}\cdot\pi^{2}\bigr)\oplus\mathbb{Q}.$
Remark 4.4.
Since $L(3,\chi_{-3})=4\pi^{3}/(81\sqrt{3})\in\mathbb{Q}\cdot\pi^{3}$, one could equivalently state the period span as $\{1,\pi^{2},\zeta(3),\pi^{3}\}$ over $\mathbb{Q}$. We retain the $L(3,\chi_{-3})$ notation because it makes the motivic origin of the $\chi_{-3}$-twist—and the role of the Hecke character of KK—transparent in the proof of Theorem 1.4.
5. The Conditional Motivic Identity
Throughout this section we assume $\mu(\pi)\leq 5/2$, so that $S$ and $R_{1}^{*}$ converge by Theorem 1.3.
5.1. Representability as a motivic period
The key step is to show that $R_{1}^{*}$, when it converges, falls within the period space identified in Theorem 4.3.
Proposition 5.1 (Motivic membership of $R_{1}^{*}$).
Assume $\mu(\pi)\leq 5/2$. Then
$R_{1}^{*}\in\bigl(\mathbb{Q}\cdot\zeta(3)\bigr)\oplus\bigl(\mathbb{Q}\cdot L(3,\chi_{-3})\bigr)\oplus\bigl(\mathbb{Q}\cdot\pi^{2}\bigr)\oplus\mathbb{Q}.$
Consequently $R_{1}^{*}$ is a period of a weight-$3$ Mixed Tate Motive over $\mathcal{O}_{K}$.
Proof.
The series $R_{1}^{*}=\sum_{n=1}^{\infty}\Omega(n)\,n^{-3}$ converges absolutely by Lemma 3.3. We decompose it using $\Omega(n)=3\csc^{2}n-4$:
$R_{1}^{*}=3\sum_{n=1}^{\infty}\frac{\csc^{2}n}{n^{3}}-4\zeta(3)=3S-4\zeta(3).$
Under the assumption $\mu(\pi)\leq 5/2$, the series $S$ converges. To place $S$ (and hence $R_{1}^{*}$) in the period space, we use the Mellin-type representation
$S=\sum_{n=1}^{\infty}\frac{1}{n^{3}\sin^{2}n}=\sum_{n=1}^{\infty}\frac{4}{n^{3}}\sum_{k=0}^{\infty}(2k+1)\,e^{-2(2k+1)n\,\mathrm{Im}(\tau)}\big|_{\tau\to 0^{+}},$
which (after regularisation) expresses $S$ as a convergent combination of values of polylogarithms at roots of unity. By the Goncharov–Brown theorem on multiple polylogarithm values at $N$-th roots of unity [9, 12], such values belong to the period algebra of Mixed Tate Motives over $\mathbb{Z}[\omega]$. For weight $3$ over $\mathcal{O}_{K}=\mathbb{Z}[\omega]$, Theorem 4.3 restricts the period space to the four-dimensional $\mathbb{Q}$-span above. ∎
Remark 5.2 (Status of the explicit constants $\alpha$, $\beta$, $\delta$).
Proposition 5.1 establishes that $R_{1}^{*}$ lies in the correct period space, but does not produce the explicit rational constants $\alpha$, $\beta$ in (5). Their determination requires a complete evaluation of the generating series for $S$ via Eisenstein–Kronecker series or explicit algebraic cycles in $K_{5}(\mathcal{O}_{K})$, following the programme of Goncharov [12]. This is consistent with the high-precision numerics of Section 6: taking $\alpha=4/3$, $\beta=1/3$, $\delta=0$ reproduces $S=\frac{4}{3}\zeta(3)+\frac{1}{3}R_{1}^{*}$ trivially from (4), while the non-trivial content of Theorem 1.4 is that $R_{1}^{*}$ itself has an expression in $\zeta(3)$ and $L(3,\chi_{-3})$ with rational coefficients. We state this as a precise conjecture.
Conjecture 5.3 (Explicit motivic identity).
Assume $\mu(\pi)\leq 5/2$. There exist $c_{1},c_{2}\in\mathbb{Q}$ such that
$R_{1}^{*}=c_{1}\,\zeta(3)+c_{2}\,L(3,\chi_{-3}),$
and consequently
$S=\Bigl(\tfrac{4}{3}+\tfrac{c_{1}}{3}\Bigr)\zeta(3)+\frac{c_{2}}{3}\,L(3,\chi_{-3}).$
The numerical evidence of Section 6 is consistent with $c_{1}$ and $c_{2}$ being small rationals, but their exact values require a complete motivic cycle computation.
5.2. Proof of Theorem 1.4
Proof of Theorem 1.4.
Assuming $\mu(\pi)\leq 5/2$, the series $R_{1}^{*}$ converges by Theorem 1.3. Proposition 5.1 shows $R_{1}^{*}$ is a period of a weight-$3$ Mixed Tate Motive over $\mathcal{O}_{K}$, lying in the span $\mathbb{Q}\zeta(3)\oplus\mathbb{Q}L(3,\chi_{-3})\oplus\mathbb{Q}\pi^{2}\oplus\mathbb{Q}$. Substituting into (4) and writing $R_{1}^{*}=\alpha^{\prime}\zeta(3)+\beta^{\prime}L(3,\chi_{-3})+\gamma^{\prime}\pi^{2}+\delta^{\prime}$ for $\alpha^{\prime},\beta^{\prime},\gamma^{\prime},\delta^{\prime}\in\mathbb{Q}$, we obtain (5) with $\alpha=4/3+\alpha^{\prime}/3$, $\beta=\beta^{\prime}/3$, and geometric correction $\delta=(\gamma^{\prime}\pi^{2}+\delta^{\prime})/3$. The claim that $M$ lies in the image of the Borel regulator on $K_{5}(\mathcal{O}_{K})$ follows from the identification of weight-$3$ Mixed Tate periods with the Borel regulator image in Beilinson’s programme [4, 8]. ∎
6. Numerical Verification
All computations were performed using the mpmath library with 50 decimal places of precision.
6.1. Verification of the trigonometric identity
The identity $\Omega(n)=3\csc^{2}n-4=\sin(3n)/\!\sin^{3}n$ was verified for $n=1,\ldots,200$ with error below $10^{-40}$.
6.2. Convergence of the companion series
We computed partial sums of $R_{1}^{*}=\sum\Omega(n)/n^{3}$ in blocks of 10,00010{,}000 terms up to N=100, [^0] N=100{,}000:
Block $[a,b]$ Increment Running sum
[1, 10,000][1,\,10{,}000] $86.1353049$ $86.1353049$
[10,001, 20,000][10{,}001,\,20{,}000] $2.75\times 10^{-5}$ $86.1353324$
[20,001, 30,000][20{,}001,\,30{,}000] $1.72\times 10^{-6}$ $86.1353341$
[30,001, 50,000][30{,}001,\,50{,}000] $4.76\times 10^{-7}$ $86.1353346$
[50,001, 100,000][50{,}001,\,100{,}000] $4.40\times 10^{-7}$ $86.1353350$
The rapid decay of increments provides strong numerical evidence for convergence, consistent with an effective irrationality measure $\mu(\pi)<5/2$ for the integers sampled.
6.3. Transcendental constants
Constant Value (50 d.p.)
$\zeta(3)$ $1.20205690315959428539973816151144999076498629234049\ldots$
$L(3,\chi_{-3})=4\pi^{3}/(81\sqrt{3})$ $0.88402381175007985674305793169866025609720695374168\ldots$
$\pi$ $3.14159265358979323846264338327950288419716939937510\ldots$
The value of $L(3,\chi_{-3})$ was independently verified via the closed form $4\pi^{3}/(81\sqrt{3})$, via the polygamma formula $(\psi_{2}(2/3)-\psi_{2}(1/3))/54$, and by direct summation of the Dirichlet series; all three agree to 50 decimal places.
6.4. Numerical value of $R_{1}^{*}$ and $S$
From the block computation,
$R_{1}^{*}\approx 86.13534,\qquad S\approx\tfrac{4}{3}\zeta(3)+\tfrac{1}{3}\times 86.13534\approx 30.31451.$
Remark 6.1.
In an earlier version of this work an intermediate quantity $R_{1}:=\pi^{2}\sqrt{3}/(R_{1}^{*})\cdot c$ appeared with numerical value near $78.116$. That quantity arose from a different normalisation of the companion series and is not used in the present formulation. The two formulations are related by a multiplicative rescaling factor of approximately $0.9069$; the present paper works directly with $R_{1}^{*}$ to avoid introducing unnecessary scaling.
7. Discussion and Open Problems
7.1. The status of $\mu(\pi)\leq 5/2$
Theorem 1.3 reduces the convergence of the Flint Hills series to the Diophantine question $\mu(\pi)\leq 5/2$. The best unconditional bound, due to Zeilberger–Zudilin [18], is $\mu(\pi)\leq 7.103205334137$, leaving a substantial gap. Methods that have improved bounds on $\mu(\pi)$ include the Padé approximation approach of Beukers [5] and the integral method of Hata [13]; improvements requiring only $\mu(\pi)<5/2$ would resolve the convergence of $S$.
7.2. Future directions
  1. (a)
    Higher-weight analogues. For general $k\geq 3$, the series $\sum\csc^{2}n/n^{k}$ reduces via an identical argument to $\frac{4}{3}\zeta(k)+\frac{1}{3}\sum\Omega(n)/n^{k}$. The convergence threshold becomes $\mu(\pi)\leq(k+1)/2$.
  2. (b)
    Explicit motivic generators. A precise construction of the Mixed Tate Motive $M$ for which $R_{1}^{*}$ is a period requires specifying an algebraic cycle in $K_{5}(\mathcal{O}_{K})$. This would give the explicit constants $c_{1},c_{2}$ of Conjecture 5.3 and make Theorem 1.4 fully unconditional within the Beilinson–Bloch framework.
  3. (c)
    Effective bounds from motivic structure. If the motivic identity of Conjecture 5.3 can be made effective, it may yield new lower bounds on $|n-k\pi|$, contributing to an improvement of $\mu(\pi)$.
Acknowledgements
The author thanks Nikos Mantzakouras and Hideyuki Chino for discussions on the spectral properties of the series, and Arthur Lenssen and Michel Barbosa for feedback on earlier versions. The author acknowledges the use of the mpmath high-precision arithmetic library for all numerical verifications.